OpenStudy (luigi0210):

Differentiate:

OpenStudy (luigi0210):

\[\LARGE y=2^{cot~x}\]

OpenStudy (agent0smith):

Use logs \[\Large \ln y = \cot x \ln 2\]Now use implicit differentiation

OpenStudy (luigi0210):

So the \[\LARGE lny\] becomes \[\LARGE \frac{1}{y'}\] right? For the right I use product rule?

OpenStudy (anonymous):

No the \(\bf ln(y)\) becomes:\[\bf \frac{ d }{ dx }\ln(y)=\frac{ 1 }{ y }*\frac{ dy }{ dx }\] This is a result of the chain rule.

OpenStudy (agent0smith):

For the right you have a constant (ln2) multiplied by cotx, no need for product rule.

OpenStudy (luigi0210):

Sorry, kind of re-learning everything again.

OpenStudy (agent0smith):

A somewhat easy way of remembering derivative of ln\[\LARGE \frac{ d }{ dx } \ln f(x) = \frac{ f'(x) }{ f(x) }\]

OpenStudy (shamil98):

and \[(\cot x)' = - \csc^2x = \frac{ -1 }{ \sin^2x } = -(1+\cot^2x)\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx }= 2^{cotx}.\ln2(-cosec^2x)\]

OpenStudy (shamil98):

\[\huge y = 2^{\cot x}\] \[\huge \ln y= \ln 2 *\cot x\] \[\huge \frac{ 1 }{ y }y' = -\csc^2x\] get y' by itself.. \[\huge y' = y(-\csc^2x)\] replace y with the original equation \[\huge y' = 2^{\cot x} * (-\csc^2x)\]

OpenStudy (anonymous):

@shamil98 You forgot the ln(2) on the RHS

OpenStudy (shamil98):

\[\huge (\ln(x))' = \frac{ 1 }{ x }\]

OpenStudy (shamil98):

oh my bad , \[\huge y ' = 2^{\cot x} * (\ln 2 * -\csc^2x)\]

OpenStudy (luigi0210):

Alright, I see.. thanks guys!